# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v1

@@ 4.5.2 Image Consistency Loss
In addition to the common likelihood loss evaluated in the latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
@@
where $\omega_{1},\omega_{2}$ are adjustable hyperparameters to balance the loss. An overview of DPBridge’s training pipeline shown in Algorithm.1.
@@
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
2:repeat
3:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
4:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
5:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:     $\mathbf{z}_{t}=n_{t}\mathbf{z}_{T}+m_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
7:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
8:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
9:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
10:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
11:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
12:until converged
@@
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
2:$\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
3:for $t=T,\dots,1$ do
4:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
5:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
6:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
7:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
8:end for
9:$\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
10:return $\mathbf{y}$
@@ 4.6 Distribution Alignment Normalization
A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel. This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer. To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions. Specifically, we normalize the intermediate latents computed from Eqn.(3) to match the form of a standard diffusion process:
@@
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
4.7 Network Architecture
To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate and terminal state samples along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [32], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
@@ 4.8 Accelerate Inference
Similar to the idea of DDIM [52], the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path. Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained. The accelerated inference process can be thus described as:
$\begin{split}\mathbf{z}_{t_{n}}&=n_{t_{n}}\mathbf{z}_{T}+m_{t_{n}}\hat{\mathbf{z}}_{0}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-n_{t_{n+1}}\mathbf{z}_{T}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process. We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Eqn.(5), the process becomes a Markovian bridge. The basic inference process of DPBridge is summarized in Algorithm.2. If adopt acceleration, one can utilize Eqn.(13) during implementation.
@@ 5 Experiments
We evaluate the proposed DPBridge approach mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction. We’ll start by a brief introduction of experiment settings of these two tasks, followed by quantitative and qualitative analysis. Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [45] and Virtual KITTI [8], which cover both indoor and outdoor scenes, for training. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [51], KITTI [22], ScanNet [15], ETH3D [49] and DIODE [56]. For surface normal prediction, we additionally incorporate iBims [33] and Sintel [7] datasets in the evaluation pipeline. For fair comparison, we use the same validation split as in previous methods [32].
@@ 5.2 Implementation Details
We implement DPBridge based on Stable Diffusion (SD) v1.5 [46] as backbone. During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 2 days to complete on a single Nvidia RTX 3090Ti GPU.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
@@
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
@@ 5.3 Quantitative and Qualitative Results
Tab.1 and Tab.2 summarize the quantitative metrics of DPBridge compared with diffusion-based and bridge-based counterparts, who are most closely related to our work. The reason we use SDv1.5 instead of more advanced SDv2.1 as backbone is that our current pipeline does not support v-prediction parameterization as in SDv2.1, and this incompatibility will undermine the model’s knowledge transfer ability. We’ll present more detail analysis in the ablation study. For a fair comparison, we re-implement all baseline methods using the same SDv1.5 backbone as we do, and report the results. We also conduct comparison with feed-forward and diffusion-based approaches based on SDv2.1 backbone, and the complete results is presented in the Appendix. DPBridge outperforms most methods in our evaluation setting and even be comparable with state-of-the-art that is trained with SDv2.1. Specifically, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge process indeed benefits the perception performance. In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
Figure 3: Qualitative comparison of monocular depth estimation methods on KITTI and ETH3D dataset. DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
Figure 4: Qualitative comparison between our (above) and Marigold trained on SDv1.5 (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
The qualitative results for depth estimation is conducted on KITTI and ETH3D datasets, as shown in Fig. 3. We can see that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range. In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows. We also present visualizations on in-the-wild images for surface normal prediction task. DPBridge is able to predict high-quality surface normals equipped with fine-grained textures. This strong generalization capability can be attribute to that we effectively harness extensive visual prior knowledge inherent in the pretrained diffusion backbone.
Method NYUv2 ScanNet iBims-1 Sintel
[…]
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge (Ours) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
@@
26.8 85.7 33.6 57.7
8.8 93.0 14.0 83.1
8.5 93.3 13.4 83.9
7.5 94.9 13.1 84.6
6.9 95.7 11.2 87.3
Table 3: The ablation study regarding the design components of DPBridge. FT stands for utilizing pretrained diffusion backbone during training. DAN denotes distribution aligned normalization technique. IC indicates image consistency loss.
Parameterization Backbone Steps NYUv2 KITTI
[…]
50 8.2 94.5 12.7 85.0
50 23.7 59.1 30.5 56.7
50 7.6 95.6 12.0 85.8
50 7.7 94.6 12.4 85.2
50 22.8 62.4 28.5 60.1
5 12.1 86.5 19.4 66.5
10 7.7 95.1 12.5 84.1
50 6.9 95.7 11.2 87.3
100 6.8 95.1 11.4 87.5
200 6.6 95.6 12.3 84.8
Table 4: The ablation study regarding the hyperparameter choice of diffusion bridge process. Options include different network parameterizations, the version of SD to use as backbone, and the number of denoising steps during inference.

### v2

@@ 4.4.2 Image Consistency Loss
In addition to the likelihood loss evaluated in latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
@@
where $\omega_{1},\omega_{2}$ are hyperparameters to balance the loss.
@@
0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
1:  repeat
2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
3:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
4:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
5:     $\mathbf{z}_{t}=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}$
6:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
7:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
8:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
9:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
10:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
11:  until Converged
@@
0:  Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
2:  for $t=T,\dots,1$ do
3:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
4:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
5:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
6:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
7:  end for
8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
9:  return $\mathbf{y}$
@@ 4.5 Distribution Alignment Normalization
A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel. This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer. To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions. Specifically, we normalize the intermediate latents computed from Equation 3 to match the form of a standard diffusion process:
@@
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone. An overview of DPBridge’s training pipeline shown in Algorithm 1.
@@ 4.6 Accelerate Inference
Similar to the idea of DDIM (Song et al., 2020a), the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path. Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained. The accelerated inference process can be thus described as:
$\begin{split}\mathbf{z}_{t_{n}}&=m_{t_{n}}\hat{\mathbf{z}}_{0}+n_{t_{n}}\mathbf{z}_{T}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}-n_{t_{n+1}}\mathbf{z}_{T}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process. We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Equation 5, the process becomes a Markovian bridge. The basic inference process of DPBridge is summarized in Algorithm 2. If adopt acceleration, one can utilize Equation 13 during implementation.
@@ 5 Experiments
We conduct evaluation mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction. We’ll start by a brief introduction of experiment settings as well as implementation details, followed by quantitative and qualitative analysis. Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim (Roberts et al., 2021) and Virtual KITTI (Cabon et al., 2020), which cover both indoor and outdoor scenes, for training. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 (Silberman et al., 2012), KITTI (Geiger et al., 2012), ScanNet (Dai et al., 2017), ETH3D (Schops et al., 2017) and DIODE (Vasiljevic et al., 2019). For surface normal prediction, we additionally incorporate iBims (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets. For fair comparison, we use the same validation split as in previous methods (Ke et al., 2024).
@@ 5.2 Implementation Details
We implement DPBridge based on Stable Diffusion v1.5 (SD1.5) and v2.1-base (SD2.1) (Rombach et al., 2022) and report their results respectively to shown compatibility with different backbones. During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE Avg.
@@
DiverseDepth (Yin et al., 2020) 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.4
MiDaS (Ranftl et al., 2020) 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
LeReS (Yin et al., 2021) 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
Omnidata (Eftekhar et al., 2021) 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
HDN (Zhang et al., 2022) 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
DPT (Ranftl et al., 2021) 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 6.9
Depth Anything V2 (Yang et al., 2024) 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
Marigold (Ke et al., 2024) 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 5.8
DepthFM (Gui et al., 2024) 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
Geowizard (Fu et al., 2025) 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.1
E2E-FT (Garcia et al., 2024) 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.4
I2SB (Liu et al., 2023) 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1 13.9
BBDM (Li et al., 2023) 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7 14.1
DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 6.1
DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.6
Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
Method Model Type NYUv2 ScanNet iBims-1 Sintel Avg.
[…]
EESNU (Bae et al., 2021) 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.5
Omnidata v1 (Eftekhar et al., 2021) 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.1
Omnidata v2 (Kar et al., 2022) 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 5.6
DSINE (Bae & Davison, 2024) 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.2
Marigold (Ke et al., 2024) 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 5.8
Geowizard (Fu et al., 2025) 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.1
E2E-FT (Garcia et al., 2024) 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 7.3
DPBridge (SD2.1) 14.8 65.3 13.4 68.5 17.1 68.0 25.0 46.3 1.4
Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@ 5.3 Quantitative and Qualitative Results
Table 1 and Table 2 summarize the quantitative metrics of DPBridge compared with state-of-the-art (SOTA) methods from both discriminative and generative fields, all in zero-shot manner. In depth estimation, DPBridge consistently matches SOTA results, while demonstrating robust generalization across datasets. It achieves 4.4 AbsRel, 97.2 % $\delta_{1}$ on NYUv2, and 4.5 AbsRel, 97.4% $\delta_{1}$ on ScanNet. Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 14.8 and $11.25^{\circ}$ accuracy of $65.3\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 17.1, $11.25^{\circ}:68.0\%$) and Sintel (mean: 25.0, $11.25^{\circ}:46.3\%$). These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks. Notably, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge construction indeed benefits the perception performance. In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
We further provide qualitative examples for both tasks in comparison with SOTA diffusion-based methods (Ke et al., 2024; Fu et al., 2025; Garcia et al., 2024), which are most closely related to our work. Figure 3 shows that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range. In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows. Figure 4 presents visualizations on in-the-wild images for surface normal prediction task, where we can see that DPBridge is able to predict high-quality surface normals equipped with fine-grained textures.
Figure 3: Qualitative comparison of diffusion-based depth estimation methods on KITTI and ETH3D dataset. DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
Figure 4: Qualitative comparison between DPBridge (above) and Marigold (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
@@
21.8 83.7 30.6 62.7
6.4 94.0 11.2 87.0
6.2 95.8 11.0 87.5
5.0 96.3 10.3 90.1
4.4 97.6 9.8 91.6
Table 3: Ablation studies regarding different design components. FT stands for finetuning pretrained backbone. DAN denotes distribution aligned normalization technique. IC indicates image consistency loss.
Figure 5: Ablation of sampling steps. The performance improves as the number of sampling steps increases, while we observe saturation after 10 steps.